# 22-large sets are sets of Bohr recurrence

### v2

@@ 1 Introduction
If a set $S$ is not a set of Bohr recurrence, then it is readily seen not to be large. One can color $n$ by taking the map $n\to(\{\alpha_{1}n\},\cdots,\{\alpha_{d}n\})\in(\mathbb{R}/\mathbb{Z})^{d}$, partitioning $(\mathbb{R}/\mathbb{Z})^{d}$ into $\varepsilon$-balls, and coloring $n$ by which $\varepsilon$-ball its image lands in. More generally, if $S$ is not a set of “nil-Bohr recurrence”, then it is not large. We are ready to state our main result.
@@
While we will not formally define “nil-Bohr recurrence” in the paper, essentially it is the higher-degree generalization of Bohr recurrence, so for instance if $s^{2}$ is far from a multiple of $\alpha$ for $s\in S$ then the identity $(x+2s)^{2}-2(x+s)^{2}+x^{2}=2s^{2}$ gives an obstruction to largeness. One can replace $s^{2}$ with any higher degree polynomial or “generalized polynomial” involving floor functions, see [1] for details.
@@ 3 The Proof
Hence, there are only $(N_{1}/N_{0})^{O_{d,1/\delta}(1)}=k^{O_{d,1/\delta}(1)}$ possible orbits, and in a random coloring each one has a chance exponentially small in $k$ of being monochromatic. So if $k$ is chosen large enough, then we will not have length $N_{2}$ arithmetic progressions that are monochromatic.
@@ 4 Further Directions
In general, a set is not “large” in the sense of [3] if it is not a set of (topological) multiple recurrence for some dynamical system. Here we have handled torus rotations, which are a special kind of dynamical system. One could ask also about increasingly more general classes of dynamical systems such as nilsystems, skew products, and general distal systems. Since every dynamical system can be built from distal systems and weak mixing systems by the Furstenberg-Zimmer structure theorem ([7] [13]), this is perhaps a potential approach to settling the $2$-large versus large conjecture in the positive.

### v3

@@ 1 Introduction
If a set $S$ is not a set of Bohr recurrence, then it is readily seen not to be large. One can color $n$ by taking the map $n\to(\{\alpha_{1}n\},\cdots,\{\alpha_{d}n\})\in(\mathbb{R}/\mathbb{Z})^{d}$, partitioning $(\mathbb{R}/\mathbb{Z})^{d}$ into $\varepsilon$-balls, and coloring $n$ by which $\varepsilon$-ball its image lands in. More generally, if $S$ is not a set of “nil-Bohr recurrence”, then it is not large. We are ready to state our main result, which resolves a conjecture from [6].
@@
While we will not formally define “nil-Bohr recurrence” in the paper. It is the higher-degree generalization of Bohr recurrence, so for instance if $s^{2}$ is far from a multiple of $\alpha$ for $s\in S$ then the identity $(x+2s)^{2}-2(x+s)^{2}+x^{2}=2s^{2}$ gives an obstruction to largeness. One can replace $s^{2}$ with any higher degree polynomial or “generalized polynomial” involving floor functions. See [1] for details.
@@ 3 The Proof
Hence, there are only $(N_{1}/N_{0})^{O_{d,1/\delta}(1)}=k^{O_{d,1/\delta}(1)}$ possible orbits, and in a random coloring each one has only an exponentially small chance $2^{-\Omega(k)}$ of being monochromatic. So if $k$ is chosen large enough, then we will not have length $N_{2}$ arithmetic progressions that are monochromatic.
@@ 4 Further Directions
In general, a set is not “large” in the sense of [3] if it is not a set of (topological) multiple recurrence for some dynamical system. Here we have handled torus rotations, which are a special kind of dynamical system. One could ask also about increasingly more general classes of dynamical systems such as nilsystems, skew products, and general distal systems. A potential approach to settling the $2$-large versus large conjecture in the positive is to settle it for some classes of dynamical systems and then to use an appropriate structure theorem to extend the result to all topological dynamical systems. A similar approach to Katznelson’s Question is outlined in [7].